# Noncooperative Equilibrium Selection via a Trading-based Auction

### v2

@@ Abstract
Noncooperative multi-agent systems often face coordination challenges due to conflicting preferences among agents. In particular, agents acting in their own self-interest can settle on different equilibria, leading to suboptimal outcomes or even safety concerns. We propose an algorithm named trading auction for consensus (TACo), a decentralized approach that enables noncooperative agents to reach consensus without communicating directly or disclosing private valuations. TACo facilitates coordination through a structured trading-based auction, where agents iteratively select choices of interest and provably reach an agreement within an a priori bounded number of steps. A series of numerical experiments validate that the termination guarantees of TACo hold in practice, and show that TACo achieves a median performance that minimizes the total cost across all agents, while allocating resources significantly more fairly than baseline approaches.
@@
Decentralized control, Equilibrium selection, Noncooperative games, Multi-agent systems
@@ 1 Introduction
Multi-agent systems often require consensus among agents, yet such consensus can be challenging due to conflicts and disagreements. These conflicts stem from differences in the strategic priorities of the agents, making coordination difficult without communication or external mediation. A key example is the equilibrium selection problem [1], where agents must choose one outcome from multiple Nash equilibria—stable outcomes where no agent can unilaterally improve their payoff. Without a mechanism to enforce a consensus, these disagreements can cause suboptimal outcomes and even safety risks in critical applications [2, 3].
Figure 1: Illustration of a waypoint merging scenario in air traffic management, where multiple aircraft must coordinate their estimated time of arrival (ETAs) at a shared waypoint. Each equilibrium represents a possible sequence of arrival times that ensures safe time separation between aircraft, highlighting the need for consensus to avoid conflicts and inefficiencies.
@@
The contributions of this work are threefold: (i) We introduce the trading auction for consensus (TACo) algorithm, which enables consensus in noncooperative settings without requiring a centralized coordinator or direct communication, and while preserving privacy. (ii) We prove that TACo enforces an agreement in finite time, and derive an explicit upper bound on the number of steps required for all agents to reach this mutual agreement. (iii) We validate our theoretical findings through numercial experiments and demonstrate that TACo outperforms existing negotiation methods regarding social optimality and fairness in a practical equilibrium selection scenario.
@@ 2 Related Work
A similar trend appears in addressing the equilibrium selection problem, where agents must coordinate among multiple conflicting outcomes. Many approaches rely on a centralized coordinator, requiring agents to share private information or follow the coordinator’s directives [4, 5, 6, 13, 14, 7, 15, 8, 9]. Others assume cooperativeness or shared beliefs, which bypass individual preferences [10, 11, 1, 12]. While effective in achieving a desirable equilibrium, these approaches are less suited for fully noncooperative systems due to their reliance on direct communication and lack of procedural rationality, where agents have incentives to participate based on self-interest.
Several works tackle equilibrium selection in noncooperative systems [12, 1, 29, 10, 11, 30, 31]. These include enforcing unique equilibria by modifying utility functions [29] or employing structured processes like the Linear Tracing Procedure, which relies on shared initial beliefs [10, 11]. While ensuring rationality during equilibrium seeking, these methods assume a coordinator or shared preferences during equilibrium selection.
Negotiation algorithms provide another approach to decentralized consensus among noncooperative agents [16]. Common methods include voting, bargaining, and auction-based mechanisms. Voting is widely used but often fails to ensure rationality, meaning outcomes may not be acceptable to all agents [16]. Bargaining relies on direct communication and is often problem-specific [32, 33, 34], limiting its applicability in general decentralized negotiation environments. In contrast, auctions align agents’ incentives with self-interest while preserving privacy (i.e., agents need not disclose sensitive private information) [35, 32, 36, 16, 37]. However, auctions primarily address resource allocation rather than single-choice problems like equilibrium selection, highlighting the need for research that bridges this gap by adapting auction-based methods for decentralized, noncooperative equilibrium selection.
@@ 3 Coordination Under Conflicting Preferences
A prominent instance of conflicting choices in multi-agent systems is the equilibrium selection problem. Before explaining the problem, we first introduce the notation and framework used to model the interactions between agents.
3.1 Equilibrium selection problem
We model the interaction between agents as a noncooperative game with $n\in\mathbb{Z}_{\geq 0}$ agents, where each agent $i\in[n]\equiv\{1,2,\ldots,n\}$ selects an action $x_{i}$ from an action set $\mathbb{X}_{i}$. The collective actions of all agents are represented by the joint action profile $\mathbf{x}=\{x_{1},x_{2},\ldots,x_{n}\}$. Each agent $i$ wishes to minimize a cost function $c_{i}(x_{i},\mathbf{x}_{\neg i})$ which depends on its own action $x_{i}$ and those of other agents, where $\mathbf{x}_{\neg i}=\mathbf{x}\setminus x_{i}$. The collection of costs of all agents is denoted as $\mathbf{c}=\{c_{1},c_{2},\ldots,c_{n}\}$.
@@ 3.1 Equilibrium selection problem
At a generalized Nash equilibrium, each agent’s strategy is optimal given the strategies of all other agents and satisfies the feasibility constraints imposed by the other agents’ actions.
@@
Example 1 (2-agent quadratic game).
Consider two agents, $i$ and $j$, aiming to minimize a shared cost while satisfying a constraint. Each agent’s action $x_{i}\in\mathbb{R}$ is chosen to solve:
$\begin{array}[]{ll}\underset{x_{i}}{\text{minimize}}&c_{i}(x_{i},x_{j})=-(x_{i}+x_{j}-1)^{2},\\ \text{subject to}&x_{i}^{2}+x_{j}^{2}\leq 1.\end{array}$ (2)
A Nash equilibrium $(x_{1}^{\star},x_{2}^{\star})$ satisfies the Karush-Kuhn-Tucker (KKT)conditions. The Lagrangian is:
$\mathcal{L}(x_{1},x_{2},\mu)=-(x_{1}+x_{2}-1)^{2}+\mu(x_{1}^{2}+x_{2}^{2}-1),$ (3)
where $\mu\geq 0$ is the Lagrange multiplier. The first-order conditions are:
$\textstyle{\frac{\partial\mathcal{L}}{\partial x_{1}}=-2(x_{1}+x_{2}-1)+2\mu x_{1}=0,}$ (4a)
$\textstyle{\frac{\partial\mathcal{L}}{\partial x_{2}}=-2(x_{1}+x_{2}-1)+2\mu x_{2}=0,}$ (4b)
$\mu(x_{1}^{2}+x_{2}^{2}-1)=0,\quad\mu\geq 0,\quad x_{1}^{2}+x_{2}^{2}\leq 1.$ (4c)
From the first two equations, we obtain:
${\mu(x_{1}-x_{2})=0,\quad x_{1}+x_{2}=\frac{2}{2-\mu}.}$ (5)
This yields two cases: i) $\mu=0$: The constraint is inactive, The feasible solutions satisfying $x_{1}^{2}+x_{2}^{2}\leq 1$ are all points on the line segment $x_{1}+x_{2}=1$ where $0\leq x_{1},x_{2}\leq 1$. ii) $\mu>0$: The constraint is active, so $x_{1}^{2}+x_{2}^{2}=1$. Since $\mu(x_{1}-x_{2})=0$, we must have $x_{1}=x_{2}$. Solving $12x_{1}^{2}=$ 1 gives $x_{1}=x_{2}=\pm\frac{1}{\sqrt{2}}$.
The critical points are:
${\{(x,1-x)\mid 0\leq x\leq 1\}\cup\left\{\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right),\left(-\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}}\right)\right\}.}$ (6)
However, $\{(x,1-x)\mid 0\leq x\leq 1\}$ are local maxima, as an agent can unilaterally decrease their cost by deviating. Thus, $\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$ and $\left(-\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}}\right)$ are Nash equilibria, demonstrating the existence of multiple equilibria in this game.
In decentralized systems, achieving consensus on which equilibrium to select is nontrivial, especially in noncooperative settings where agents have different preferences over equilibria. Without coordination, there is a risk that agents may settle on different equilibria, resulting in inefficiencies or even safety risks in the system. Furthermore, agents may not be able to communicate directly or share their preferences, further complicating the equilibrium selection problem.
@@ 4.1 Key elements of TACo
    Offer matrix ($\mathcal{O}$): The offer matrix $\mathcal{O}\in\mathbb{R}^{n\times m}$ represents the units that each agent $i$ can receive if they select choice $j$. Specifically, $\mathcal{O}_{ij}$ denotes the number of units that agent $i$ will receive when they choose choice $j$, determined by the offers made by other agents.
@@
    Pay matrix ($\mathcal{P}$): The pay matrix $\mathcal{P}\in\mathbb{R}^{n\times m}$ captures the cost each agent incurs when selecting a particular choice. $\mathcal{P}_{ij}$ is the number of units that agent $i$ has to pay to others when they choose choice $j$.
@@
    Cost matrix ($\mathcal{C}$): The cost matrix $\mathcal{C}\in\mathbb{R}^{n\times m}$ reflects the intrinsic cost associated with each agent’s choice. $\mathcal{C}_{ij}$ represents the original cost experienced by agent $i$ when they select choice $j$.
@@
    This reduction adjusts the scale of assets exchanged in the auction, allowing for a controlled decrease in trading volume over time or under certain conditions, which can influence the convergence behavior of the auction process.
@@
    $\mathcal{J}=\text{diag}(\mathbf{b})\cdot(\mathcal{O}-\mathcal{P})-\mathcal{C}.$ (8)
    This equation accounts for each agent’s private valuation b, the offers received $\mathcal{O}$, the payments made $\mathcal{P}$, and the intrinsic costs $\mathcal{C}$, resulting in the total profit for each possible choice.
@@ 4.2 Auction mechanism
A cycle, $\sigma$, is a finite sequence of agent-profit matrix tuples,
@@
in which a tuple $(i^{k_{\ell+1}},\mathcal{J}^{k_{\ell+1}})$ revisits any previously encountered tuple, i.e.,
$(i^{k_{\ell+1}},\mathcal{J}^{k_{\ell+1}})=(i^{k_{1}},\mathcal{J}^{k_{1}}).$
Here, $\ell$ is the length of the cycle.
@@
1:$n$: number of agents, $m$: number of choices, $\epsilon$: termination tolerance
2:$\mathcal{C}\in\mathbb{R}^{n\times m}$, $d_{0}\in\mathbb{R}_{>0}$, $\gamma\in(0,1)$
3:Initialize offer matrix $\mathcal{O}\in\mathbb{N}^{n\times m}$ to $\mathbf{0}_{n\times m}$
4:Initialize pay matrix $\mathcal{P}\in\mathbb{N}^{n\times m}$ to $\mathbf{0}_{n\times m}$
@@
10:     Choose a playing agent $i$ in a sequential manner.
@@
15:     $j^{\star}\leftarrow\underset{j\in[m]}{\arg\max}(\mathcal{J}_{ij})$triangleright Agent $i$ selects the choice $j^{\star}$ that maximizes its profit
@@
19:     $\text{{selections}}_{i}\leftarrow j^{\star}$triangleright Record agent $i$’s choice
@@
24:              $\max_{j}\mathcal{J}_{ij}-\min_{j}\mathcal{J}_{ij}<\epsilon,\,\forall j\in\text{cycle}$triangleright Check termination condition
@@
32:return $\text{mode}(\textit{selections})$triangleright Return the most frequent value in selections
@@
    Cycle Detection and Trading Unit Reduction: The trading unit $d$ is reduced by the decrement factor $\gamma$ when a cycle is detected. A cycle is identified by checking whether the same pair of profit matrix $\mathcal{J}$ and active agent at a specific step has occurred in any previous steps. Tracking $\mathcal{J}$ is equivalent to tracking $(\mathcal{O}-\mathcal{P})$, since $b_{i}$ and $C_{ij}$ are constants based on eq. 9. Agents can independently detect cycles by monitoring changes in the publicly observable matrices $\mathcal{O}$ and $\mathcal{P}$. When a cycle is detected, the trading unit is updated as $d_{r+1}\leftarrow\gamma d_{r}$, where $d_{r}$ denotes the trading unit after $r$ cycle detections, with the initial trading unit defined as $d_{0}$. Additionally, the current record of profit matrices is cleared to restart the cycle detection process.
    This reduction of the trading unit is a common practice in real bargaining scenarios, where participants tend to adjust their bids or pricing in smaller increments over time to facilitate reaching a consensus [38]. The relationship governing the trading unit after $r$ cycle detections can be expressed as:
@@
    $\epsilon$-termination: The auction continues until consensus is reached. The process terminates when the difference in profit between all agents’ choice options is within a fixed tolerance $\epsilon$, indicating indifference among the choices in the cycle.
@@
Consider a case with two agents ($n=2$) and two choices ($m=2$). The initial cost matrix is given by $\,\mathcal{C}=\begin{bmatrix}10&4\\ 7&9\end{bmatrix}$, and the agents’ private valuations for the traded asset are $\mathbf{b}=\begin{bmatrix}0.8&1.2\end{bmatrix}$. At each step, agents update their selections based on their calculated profit using the TACo rules. Table 1 shows how the auction progresses over time. In this example, Agent 1 initially selects option 2 because it provides the highest profit. In step 2, Agent 2 selects option 1, which maximizes its own profit. At each step, the agent currently playing updates its selection, and this updated selection is underlined in the table. The bold values in the profit matrix $\mathcal{J}$ highlight that the same profit-value and agent pair reappear at steps 3 and 5, signaling the occurrence of a cycle. This cycle is associated with selection 2, where all agents converge to the same option. The algorithm detects that the $\epsilon$-termination criterion is satisfied because the profit differences for all agents are zero, which is smaller than any positive $\epsilon$. Consequently, the algorithm terminates at step 5.

### v3

# Noncooperative Consensus via a Trading-based Auction

@@ Abstract
Noncooperative multi-agent systems often face coordination challenges due to conflicting preferences among agents. In particular, when agents act in their own self-interest, they may prefer different choices among multiple feasible outcomes, leading to suboptimal results or even safety concerns. We propose an algorithm named trading auction for consensus (TACo), a decentralized approach that enables noncooperative agents to reach consensus without communicating directly or disclosing private valuations. TACo facilitates coordination through a structured trading-based auction, where agents iteratively select choices of interest and provably reach an agreement within an a priori bounded number of steps. A series of numerical experiments validate that the termination guarantees of TACo hold in practice, and show that TACo achieves a median performance that minimizes the total cost across all agents, while allocating resources significantly more fairly than baseline approaches.
@@
Decentralized control, Consensus problem, Noncooperative games, Multi-agent systems
@@ 1 Introduction
Multi-agent systems often require consensus among agents, yet such consensus can be challenging due to conflicts and disagreements. These conflicts stem from differences in the strategic priorities of the agents, making coordination difficult without communication or external mediation. A representative example of such challenges arises when agents must select one shared outcome from multiple feasible choices that reflect their individual strategic preferences. Without a mechanism to enforce a consensus, these disagreements can cause suboptimal outcomes and even safety risks in critical applications [1, 2] as illustrated in Figure 1.
Figure 1: Illustration of a waypoint merging scenario in air traffic management, where multiple aircraft must coordinate their estimated times of arrival (ETAs) at a shared waypoint. Each possible outcome represents a feasible ordering of aircraft that maintains safe time separation, highlighting the need for consensus among self-interested agents to avoid conflicts.
@@
The contributions of this work are threefold: (i) We introduce the trading auction for consensus (TACo) algorithm, which enables consensus in noncooperative settings without requiring a centralized coordinator or direct communication, and while preserving privacy. (ii) We prove that TACo enforces an agreement in finite time, and derive an explicit upper bound on the number of steps required for all agents to reach this mutual agreement. (iii) We validate our theoretical findings through numerical experiments and demonstrate that TACo outperforms existing negotiation methods with respect to social optimality and fairness in a practical equilibrium selection scenario.
@@ 2 Related Work
A related stream of research considers coordination among agents when multiple feasible outcomes exist, often framed as multi-choice problems. Most of these approaches adopt a centralized or cooperative structure, allowing agents to collectively evaluate all possible outcomes and agree on one through a common decision rule [3, 4, 5, 13, 14, 6, 15, 7, 8]. While such frameworks can ensure consistency in the final outcome, they require agents to disclose private cost information or require agents to follow decision directives issued by a central authority, making them impractical in decentralized settings where privacy and autonomy are essential.
Several works tackle multi-choice problems in noncooperative systems [12, 11, 29, 9, 10, 30, 31]. Examples include enforcing unique joint decisions by adjusting agents’ utilities [29] or employing structured processes like the linear tracing procedure that rely on common initial beliefs [9, 10]. Although these methods maintain individual rationality, where agents have incentives to participate based on self-interest, they still depend on partial coordination or shared information, and thus fall short of achieving fully decentralized consensus.
Negotiation algorithms provide another approach to decentralized consensus among noncooperative agents [16]. Common methods include voting, bargaining, and auction-based mechanisms. Voting is widely used but often fails to ensure rationality, meaning outcomes may not be acceptable to all agents [16]. Bargaining relies on direct communication and is often problem-specific [32, 33, 34], limiting its applicability in general decentralized negotiation environments. In contrast, auctions align agents’ incentives with self-interest while preserving privacy (i.e., agents need not disclose sensitive private information) [35, 32, 36, 16, 37]. However, auctions primarily address resource allocation rather than single-choice problems, highlighting the need for research that bridges this gap by adapting auction-based methods for decentralized, noncooperative equilibrium selection.
@@ 3 Coordination Under Conflicting Preferences
In decentralized multi-agent systems, agents are often required to reach agreement on one outcome from a finite set of feasible alternatives, even when their individual preferences conflict. We refer to this general setting as a multi-choice consensus problem. Each agent aims to minimize its own cost or maximize its utility, but the system must converge to a single shared choice to ensure coherent operation. Such situations arise naturally in scheduling, task allocation, and traffic management, where multiple admissible joint configurations may exist but only one can be implemented in practice.
A prominent example of this class of problems is the equilibrium selection problem. In many noncooperative games, several equilibria coexist, each representing a distinct yet individually stable outcome for all agents. Although every equilibrium satisfies the best-response condition, the coexistence of multiple equilibria introduces ambiguity: agents must somehow agree on one equilibrium to realize a consistent joint outcome. This subsection illustrates this issue using standard game-theoretic notation.
3.1 Equilibrium selection as a representative multi-choice problem
We model the interaction between agents as a noncooperative game with $n\in\mathbb{Z}_{\geq 0}$ agents. Let $\mathbf{x}=\{x_{1},x_{2},\ldots,x_{n}\}$ denote the joint action profile of all agents, and $\mathbf{x}_{\neg i}=\mathbf{x}\setminus x_{i}$ the actions of all agents except $i$. Each agent $i\in[n]\equiv\{1,2,\ldots,n\}$ selects an action $x_{i}$ from a feasible action set $\mathbb{X}_{i}(\mathbf{x}_{\neg i})$. Each agent $i$ aims to minimize a cost function $c_{i}(x_{i},\mathbf{x}_{\neg i})$ that depends on its own action $x_{i}$ and those of other agents. The collection of cost functions of all agents is denoted as $\mathbf{c}=\{c_{1},c_{2},\ldots,c_{n}\}$.
@@ 3.1 Equilibrium selection as a representative multi-choice problem
At a generalized Nash equilibrium, each agent’s strategy is optimal given the strategies of all other agents and satisfies the feasibility constraints imposed by the other agents’ actions. Unless otherwise stated, all references to Nash equilibria in the remainder of this paper refer to generalized Nash equilibria.
@@
Example 1 (Two-agent quadratic game).
Consider two agents minimizing
$c_{i}(x_{i},x_{j})=-(x_{i}+x_{j}-1)^{2},\quad\text{s.t. }x_{i}^{2}+x_{j}^{2}\leq 1.$
The corresponding KKT conditions yield two distinct Nash equilibria, $(x_{1}^{\star},x_{2}^{\star})=\left(\tfrac{1}{\sqrt{2}},\tfrac{1}{\sqrt{2}}\right)$ and $(x_{1}^{\star},x_{2}^{\star})=\left(-\tfrac{1}{\sqrt{2}},-\tfrac{1}{\sqrt{2}}\right)$, demonstrating that even simple games can admit multiple feasible equilibria.
In decentralized systems, achieving consensus on which feasible outcome to select is nontrivial, especially in noncooperative settings where agents have different preferences over choices. Without coordination, there is a risk that agents may converge to different choices, resulting in inefficiencies or even safety risks. Furthermore, agents may not be able to communicate directly or share their preferences, further complicating the multi-choice consensus problem.
@@ 4.1 Key elements of TACo
    Offer matrix ($\mathcal{O}$): The offer matrix $\mathcal{O}\in\mathbb{R}^{n\times m}$ represents the units that each agent $i$ can receive if it selects choice $j$. Specifically, $\mathcal{O}_{ij}$ denotes the number of units that agent $i$ will receive when it chooses choice $j$, determined by the offers made by other agents.
@@
    Pay matrix ($\mathcal{P}$): The pay matrix $\mathcal{P}\in\mathbb{R}^{n\times m}$ captures the cost each agent incurs when selecting a particular choice. $\mathcal{P}_{ij}$ is the number of units that agent $i$ has to pay to others when it chooses choice $j$.
@@
    Cost matrix ($\mathcal{C}$): The cost matrix $\mathcal{C}\in\mathbb{R}^{n\times m}$ reflects the intrinsic cost associated with each agent’s choice. $\mathcal{C}_{ij}$ represents the original cost experienced by agent $i$ when it selects choice $j$. Each row of $\mathcal{C}$ corresponds to the private cost structure of an agent and need not be shared with others, preserving privacy in how each agent evaluates the available options.
@@
    This reduction adjusts the scale of assets exchanged in the auction, allowing for a controlled decrease in trading volume over time or under certain conditions, which—later in Section 5—plays a key role in ensuring termination. Without this refinement, the trading process could continue indefinitely with large discrete updates, failing to settle into a consensus.
@@
    $\mathcal{J}=\mathrm{diag}(\mathbf{b})\cdot(\mathcal{O}-\mathcal{P})-\mathcal{C}.$ (3)
    This equation accounts for each agent’s private valuation $\mathbf{b}$, the offers received $\mathcal{O}$, the payments made $\mathcal{P}$, and the intrinsic costs $\mathcal{C}$, resulting in the total profit for each possible choice. Here, $b_{i}$ serves as an agent-specific conversion coefficient that normalizes each agent’s intrinsic cost unit in $\mathcal{C}$ to the common unit of the secondary asset used in trading.
The parameters $d$, $\gamma$, and $\varepsilon$ are public protocol constants shared by all agents. They can be predetermined by an oversight authority or jointly agreed upon before execution, and contain no private information, preserving the decentralized and privacy-preserving nature of the algorithm.
@@ 4.2 Auction mechanism
A cycle, $\sigma$, is a finite sequence of agent–profit matrix tuples,
@@
that returns to a previously encountered tuple, i.e.,
$(i^{k_{\ell+1}},\mathcal{J}^{k_{\ell+1}})=(i^{k_{1}},\mathcal{J}^{k_{1}}),$
where the $(\ell+1)$-th tuple coincides with the first, and $\ell$ denotes the length of the cycle.
@@
1:$n$: number of agents, $m$: number of choices, $\varepsilon$: termination tolerance
[…]
3:Initialize offer matrix $\mathcal{O}\in\mathbb{{R}}^{n\times m}$ to $\mathbf{0}_{n\times m}$
4:Initialize pay matrix $\mathcal{P}\in\mathbb{{R}}^{n\times m}$ to $\mathbf{0}_{n\times m}$
@@
10:  Choose a playing agent $i$ in a cyclic order.
@@
15:  $j^{\star}\leftarrow\underset{j\in[m]}{\arg\max}(\mathcal{J}_{ij})$ $\triangleright$ Agent $i$ selects the choice $j^{\star}$ that maximizes its profit
@@
19:  $\text{{selections}}_{i}\leftarrow j^{\star}$ $\triangleright$ Record agent $i$’s choice
@@
24:     $\max_{j}\mathcal{J}_{ij}-\min_{j}\mathcal{J}_{ij}<\varepsilon,\,\forall j\in\text{cycle}$ $\triangleright$ Check termination condition
@@
[…]
33:for each agent $i\in[n]$ do
34:  $\pi_{i}\leftarrow\mathcal{O}_{ij^{\star}}-\mathcal{P}_{ij^{\star}}$ $\triangleright$ Compute one-time payment at termination
35:end for
36:return $j^{\star}$, $\{\pi_{i}\}_{i=1}^{n}$, $\mathcal{O}$, $\mathcal{P}$
@@
    Cycle detection and trading unit reduction: The trading unit $d$ is reduced by the decrement factor $\gamma$ when a cycle is detected. A cycle is identified by checking whether the same pair of profit matrix $\mathcal{J}$ and active agent at a specific step has occurred in any previous steps. Tracking $\mathcal{J}$ is equivalent to tracking $(\mathcal{O}-\mathcal{P})$, since $b_{i}$ and $C_{ij}$ are constants based on eq. 4. Agents can independently detect cycles by monitoring changes in the publicly observable matrices $\mathcal{O}$ and $\mathcal{P}$. Because the playing agent index is part of each recorded state, a cycle cannot be detected until every agent has taken at least one turn. This ensures that termination is only possible after a full cyclic pass of agent updates. When a cycle is detected, the trading unit is updated as $d_{r+1}\leftarrow\gamma d_{r}$, where $d_{r}$ denotes the trading unit after $r$ cycle detections, with the initial trading unit defined as $d_{0}$. Additionally, the current record of profit matrices is cleared to restart the cycle detection process.
    This reduction of the trading unit is a common practice in real bargaining scenarios, such as price negotiations, where parties gradually reduce their adjustment steps as they move toward an agreement [38]. The relationship governing the trading unit after $r$ cycle detections can be expressed as:
@@
    $\varepsilon$-termination: The auction continues until consensus is reached. The process terminates when the difference between an agent’s maximum and minimum profit across all choices in a cycle falls within a fixed tolerance $\varepsilon$, which reflects that the agent is effectively indifferent among the available choices within the detected cycle.
  6.  6)
    Transfer settlement: After the termination criterion is satisfied, the algorithm fixes the final choice $j^{\star}$ and settles all transfers once for $j^{\star}$. Each agent’s net transfer in asset units is $\pi_{i}=\mathcal{P}_{ij^{\star}}-\mathcal{O}_{ij^{\star}}$.
@@
Consider a case with two agents ($n=2$) and two choices ($m=2$). The initial cost matrix is given by $\mathcal{C}=\begin{bmatrix}10&4\\ 7&9\end{bmatrix}$, and the agents’ private valuations for the traded asset are $\mathbf{b}=\begin{bmatrix}0.8&1.2\end{bmatrix}$. At each step, agents update their selections based on their calculated profit using the TACo rules. Table 1 shows how the auction progresses over time. At each step, the agent currently playing updates its selection, and this updated selection is underlined in the table.
In this example, agent 1 initially selects option 2 because it provides the highest profit. In step 2, agent 2 selects option 1, maximizing its own profit given agent 1’s choice. In step 3, agent 1 again chooses option 2 based on the updated profit matrix $\mathcal{J}$. In step 4, agent 2 changes its choice and selects option 1, and finally, at step 5, agent 1 again selects option 2.
The bold values in the profit matrix $\mathcal{J}$ highlight that the same profit-value and agent pair reappear at steps 3 and 5, signaling the occurrence of a cycle. This cycle is associated with selection 2, where all agents converge to the same option. The algorithm detects that the $\varepsilon$-termination criterion is satisfied because the profit differences for all agents are zero, which is smaller than any positive $\varepsilon$. Consequently, the algorithm terminates at step 5.